### v1

@@ 6 Conclusions
This work presented the rationale, target selection, and first characterisation of the sample spectroscopically monitored with FLAMES/VLT in the framework of the Binarity at LOw Metallicity (BLOeM) ESO Large Programme. BLOeM will collect 25 epochs of spectroscopy of 929 massive stars in the low metallicity conditions of the SMC in the period from October 2023 to September 2025. The sample populates eight fields within the SMC, probing several of its environments, though limited to field stars.
The goals of the survey are to use the time-dependent RVs of all targets to derive the observed and intrinsic binary fraction as a function of stellar mass and age, derive the orbits of all identified binaries, and, through this, establish fundamental properties such as the initial mass function, star formation history, surface abundance pattern, and orbital-parameter distributions of massive stars at low metallicity. Moreover, the survey will enable the identification of unique evolved binaries such as dormant OB+BH binaries, and provide testbeds for single-star evolution models via dynamical-mass measurements of eclipsing binaries.
This study was based on 9/25 epochs thus far acquired in the framework of the BLOeM survey during the first semester, with the main goal of describing the sample selection (Sect. 2), developing a data-reduction pipeline (Sect. 3), performing a spectral classification (Sect. 4), and investigating the mass and age domains probed by the sample (Sect. 5).
@@
We highlighted a few peculiar objects in our sample: eight sources are confirmed as X-ray bright, four of which have been classified as HMXBs in the past, and one of which is a promising CWB candidate. Three objects are classified as sgB[e]/LBV-like stars, and two as Be + bloated stripped-star binary candidates. Two candidate magnetic stars, classified as Of?p stars previously, are also included in the sample.
We identify 75 stars as Oe/Be stars. Since we lack the diagnostic H line, this should be considered a lower limit, although usage of low resolution Gaia spectra which cover the H line only yielded a few more candidates, amounting to an Oe/Be fraction of relative to the non-supergiant OB sample. The Oe fraction among all O stars in the sample is 13%. However, the Be fraction is merely 10%, which is lower than reported in the literature. It is likely that we are biased against Be stars in this sample, since these tend to be redder compared to a ”standard” evolutionary track, which we used to select our sample (see Sect. 5.2). The Oe sample however is likely unbiased, and so this fraction should represent the Oe fraction in the SMC.
Finally, 74 eclipsing binaries are identified, as well as eight ellipsoidal variables (Sect. 5.3 and Appendix B).
A systematic analysis of the first batch of data for RV variability, for the various subsamples of the study, is currently underway.
@@ Acknowledgements.
- Bordier et al. (2023) Bordier, E., de Wit, W. J., Frost, A. J., et al. 2023, arXiv e-prints, arXiv:2311.06131
@@
- Gilkis et al. (2021) Gilkis, A., Shenar, T., Ramachandran, V., et al. 2021, arXiv e-prints, arXiv:2102.03102
@@
- Marín Pina et al. (2024) Marín Pina, D., Rastello, S., Gieles, M., et al. 2024, arXiv e-prints, arXiv:2404.13036
@@ Appendix A Epoch dates and spectral types
5-001 | 4687500991624952064 | OGLE SMC-SC10 24428 | 15.05 | O9 Ve [Lam16] | O9.7 Vn + B | no | LC from [Lam16] |
5-002 | 4687504324519462528 | OGLE SMC113.7 34708 | 15.94 | - | B2 IV | no | - |
@@
Ard77: Ardeberg & Maurice (1977); Azz79: Azzopardi & Vigneau (1979); Azz82: Azzopardi & Vigneau (1982); Coe15: Coe & Kirk (2015); Dac70: Dachs (1970); Eva04: Evans et al. (2004); Eva06: Evans et al. (2006); Gar87: Garmany et al. (1987); Gra12: Graus et al. (2012); Hil05: Hilditch et al. (2005); Hum83: Humphreys (1983); Hum91: Humphreys et al. (1991); Lam13: Lamb et al. (2013); Mar07: Martayan et al. (2007); Mas95: Massey et al. (1995); Mas00: Massey et al. (2000); Mas02: Massey (2002); Mas04: Massey et al. (2004); Mas09: Massey et al. (2009); Men06: Mennickent et al. (2006); Mok06: Mokiem et al. (2006); Mor03: Morrell et al. (2003); Neu10: Neugent et al. (2010); Neu18: Neugent et al. (2018); Pau12: Paul et al. (2012); Paw16: Pawlak et al. (2016); Pri87: Prinja (1987); San68: Sanduleak (1968); She13: Sheets et al. (2013); Smi97: Smith Neubig & Bruhweiler (1999); Tes87: Testor & Lortet (1987); Tes01: Testor (2001); Wal83: Walborn (1983); Wal00: Walborn et al. (2000); Wal02: Walborn et al. (2002); Zas14: Zasche et al. (2014); Zic96: Zickgraf et al. (1996);
Appendix B Cross matches with additional catalogues
B.1 ESO archive
@@
B.2 Hubble UV Legacy Library of Young Stars as Essential Standards (ULLYSES)
ULLYSES is a legacy survey of the Hubble Space Telescope (HST), which includes the acquisition of high-resolution UV spectra for 128 massive stars in the SMC777https://ullyses.stsci.edu/ullyses-targets-smc.html. The programme also includes a follow-up with the X-SHOOTER spectrograph of the VLT to obtain a visual and infrared coverage of the targets Vink et al. (2023). A cross-match of the BLOeM sample with the ULLYSES sample, using a search radius of 3”, resulted in an overlap of 43 targets. From ULLYSES and XShootU there will be broad wavelength coverage of the UV and visible spectrum of these objects that will also be used to inform the analysis of these targets.
B.3 X-ray catalogues
The SMC was extensively observed in X-rays. The largest modern X-ray observatories, XMM-Newton and Chandra, which operate in 0.2-12.0 keV range, conducted surveys of the entire SMC galaxy (Laycock et al. 2010; Sturm et al. 2013). The deep observations of individual fields, such as the SMC Wing and NGC 346 star cluster have also been performed (Nazé et al. 2004; Oskinova et al. 2013). However, despite these efforts X-ray emission of individual “normal” massive OB stars is below current detection limits. On the other hand, X-ray detections of massive stars in the SMC allow to select binary stars. Specifically, X-ray detections are excellent tracers of CWBs, some of which are significantly more X-ray bright compared to single stars (Corcoran et al. 1996; Sana et al. 2006; Oskinova 2005; Nazé et al. 2007). But best of all, X-ray detections are suited to identify high-mass X-ray binaries (HMXB), where a compact object is accreting matter of its OB-type companion. HMXBs are X-ray variable, especially so are BeXRBs where the donor stars have OBe spectral type. BeXRBs are transient X-ray sources, and may remain quiescent over long periods of time and could be detected only during outbursts.

### v2

@@ 6 Conclusions
This work presents the rationale, target selection, and first characterisation of the sample spectroscopically monitored with FLAMES/VLT in the framework of the Binarity at LOw Metallicity (BLOeM) ESO Large Programme. BLOeM will collect 25 epochs of spectroscopy of 929 massive stars in the low-metallicity conditions of the SMC in the period from October 2023 to September 2025. The sample populates eight fields within the SMC, probing several of its environments, though limited to field stars.
The goals of the survey are to use the time-dependent RVs of all targets to derive the observed and intrinsic binary fraction as a function of stellar mass and age, derive the orbits of all identified binaries, and, through this, establish fundamental properties such as the initial mass function, star formation history, surface abundance pattern, and orbital-parameter distributions of massive stars at low metallicity. Moreover, the survey will enable the identification of unique evolved binaries, such as dormant OB+BH binaries, and provide testbeds for single-star evolution models via dynamical-mass measurements of eclipsing binaries.
The present study is based on 9/25 epochs acquired thus far in the framework of the BLOeM survey during the first semester, with our main goal being to describe the sample selection (Sect. 2), develop a data-reduction pipeline (Sect. 3), perform a spectral classification (Sect. 4), and investigate the mass and age domains probed by the sample (Sect. 5).
@@
We highlight a few peculiar objects in our sample: eight sources are confirmed as X-ray bright, four of which have been classified as HMXBs in the past, and one of which is a promising CWB candidate. Three objects are classified as sgB[e]/LBV-like stars, and two as Be + bloated stripped-star binary candidates. Two candidate magnetic stars, classified as Of?p stars previously, are also included in the sample.
We identify 82stars as OBe stars. As we lack the diagnostic H line, this should be considered a lower limit, although usage of low-resolution Gaia spectra covering the H line only yields a few more candidates, amounting to an OBe fraction of relative to the non-supergiant OB sample. The Oe fraction among all O stars in the sample is 13%. However, the Be fraction is merely 11%, which is lower than reported in the literature. It is likely that we are biased against Be stars in this sample, because these tend to be redder compared to a ‘standard’ evolutionary track, which we used to select our sample (see Sect. 5.2). However, the Oe sample is likely unbiased, and so this fraction should represent the Oe fraction in the SMC. Finally, 74 eclipsing binaries are identified, as well as 8 ellipsoidal variables (Sect. 5.3 and Appendix C). We are currently carrying out a systematic analysis of the first batch of data for RV variability for the various subsamples of the study.
@@ Acknowledgements.
- Bordier et al. (2024) Bordier, E., de Wit, W. J., Frost, A. J., et al. 2024, A&A, 681, A85
@@
- Gilkis et al. (2021) Gilkis, A., Shenar, T., Ramachandran, V., et al. 2021, MNRAS, 503, 1884
@@
- Gray & Corbally (2009) Gray, R. O. & Corbally, Christopher, J. 2009, Stellar Spectral Classification (Princeton University Press)
@@
- Marín Pina et al. (2024) Marín Pina, D., Rastello, S., Gieles, M., et al. 2024, A&A, 688, L2
@@ Appendix A Epoch dates and spectral types
5-001 | 4687500991624952064 | OGLE SMC-SC10 24428 | 15.05 | O9 Ve [Lam16] | O9.7 Vn + B | no | SB2, LC from [Lam16] |
5-002 | 4687504324519462528 | OGLE SMC113.7 34708 | 15.94 | - | B2 IV | no | EB [Paw13] |
@@
Ard77: Ardeberg & Maurice (1977); Azz79: Azzopardi & Vigneau (1979); Azz82: Azzopardi & Vigneau (1982); Coe15: Coe & Kirk (2015); Eva04: Evans et al. (2004); Eva06: Evans et al. (2006); Gar87: Garmany et al. (1987); Gra12: Graus et al. (2012); Hil05: Hilditch et al. (2005); Hum83: Humphreys (1983); Hum91: Humphreys et al. (1991); Lam13: Lamb et al. (2013); Mar07: Martayan et al. (2007); Mas95: Massey et al. (1995); Mas00: Massey et al. (2000); Mas02: Massey (2002); Mas04: Massey et al. (2004); Mas09: Massey et al. (2009); Men06: Mennickent et al. (2006); Mok06: Mokiem et al. (2006); Mor03: Morrell et al. (2003); Neu10: Neugent et al. (2010); Neu18: Neugent et al. (2018); Pau12: Paul et al. (2012); Paw16: Pawlak et al. (2016); Pri87: Prinja (1987); San68: Sanduleak (1968); She13: Sheets et al. (2013); Smi97: Smith Neubig & Bruhweiler (1999); Tes87: Testor & Lortet (1987); Tes01: Testor (2001); Wal83: Walborn (1983); Wal00: Walborn et al. (2000); Wal02: Walborn et al. (2002); Zas14: Zasche et al. (2014); Zic96: Zickgraf et al. (1996);
Appendix B DSS images for nebular contamination
In order to get a better handle on which stars are affected by nebular contamination, and which objects show intrinsic emission like in the case of classical OeBe stars, we investigated wide-field Digitized Sky Survey (DSS) 2-red images. We retrieved 25’ x 25’ cutouts for each of the fields and overplotted all BLOeM sources in order to investigate their local surroundings. We inspected the spectra of all sources that show overdensitites in DSS 2-red, in particular all objects classified as emission-line stars, to better distinguish between nebular contamination and classical OeBe stars. We designated with ’neb’ all objects located inside a nebulosity visible in DSS 2-red and show narrow emission lines, mainly in the H line.
In Figs. 11 and 12 we show the DSS 2-red cutouts for each of the eight fields. Here, we overplot all BLOeM sources, in particular emission-line stars classified as OeBe stars, and mark stars that are affected by nebular contamination. Some fields, for example Field 1, 4 or 6, have large nebulosities in the fields of view and many sources are affected by nebular contamination. Other fields, like fields 5, 7 or 8, are barely or not affected at all. Few objects are classical OeBe stars and additionally show nebular contamination in their spectra.
Appendix C Cross-matches with additional catalogues
C.1 ESO archive
@@
C.2 Hubble UV Legacy Library of Young Stars as Essential Standards (ULLYSES)
ULLYSES is a legacy survey of the Hubble Space Telescope (HST), which includes the acquisition of high-resolution UV spectra for 128 massive stars in the SMC888https://ullyses.stsci.edu/ullyses-targets-smc.html. The programme also includes a follow-up with the X-SHOOTER spectrograph of the VLT to obtain a visual and infrared coverage of the targets Vink et al. (2023). A cross-match of the BLOeM sample with the ULLYSES sample, using a search radius of 3”, resulted in an overlap of 43 targets. From ULLYSES and XShootU there will be broad wavelength coverage of the UV and visible spectrum of these objects that will also be used to inform the analysis of these targets.
C.3 X-ray catalogues
The SMC was extensively observed in X-rays. The largest modern X-ray observatories, XMM-Newton and Chandra, which operate in 0.2-12.0 keV range, conducted surveys of the entire SMC galaxy (Laycock et al. 2010; Sturm et al. 2013). The deep observations of individual fields, such as the SMC Wing and NGC 346 star cluster have also been performed (Nazé et al. 2004; Oskinova et al. 2013). However, despite these efforts X-ray emission of individual ‘normal’ massive OB stars is below current detection limits. On the other hand, X-ray detections of massive stars in the SMC allow to select binary stars. Specifically, X-ray detections are excellent tracers of CWBs, some of which are significantly more X-ray bright compared to single stars (Corcoran et al. 1996; Sana et al. 2006; Oskinova 2005; Nazé et al. 2007). But best of all, X-ray detections are suited to identify high-mass X-ray binaries (HMXBs), where a compact object is accreting matter of its OB-type companion. HMXBs are X-ray variable, especially so are BeXRBs where the donor stars have OBe spectral type. BeXRBs are transient X-ray sources, and may remain quiescent over long periods of time and could be detected only during outbursts.
@@ C.5 TESS photometry
Appendix D HRDs separated by field
Figure 13 shows the HRDs of each of the eight SMC fields observed in the framework of BLOeM (shown in Fig. 2). The differences between the populations are not blatant. Generally, fields 1 – 4 contain a higher number of stars ”born as O-type” () compared to fields 5 – 8, which could be anticipated given the automated way with which the fields were selected, prioritizing fields with the most massive stars (Sect. 2). Field 5 appears to contain the oldest population among the eight fields.